A Complete Bibliography of Publications in Probability Surveys

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 November 2024 Version 1.17

Title word cross-reference

m [93]. Z^{d} [113]. ψ [40]. U [18].

-ary [93]. -statistics [18]. -weak [40].

abelian [59]. absolutely [24]. Activated [113, 148]. Addendum [53, 121]. advances [20]. against [145]. algebraic [133]. algorithms [118, 2]. allocations [64]. alternative [55]. analysis [73, 44]. Anderson [91]. ansatz [99]. applications [117, 40, 122, 140, 14, 134, 71, 19, 82]. approach [123, 110, 132]. approximation [128, 9, 127, 107, 121, 124]. approximations [38]. area [31, 53, 48]. areas [31]. ary [93]. aspects [19, 120]. associated [131]. asymmetric [101]. Asymptotic [143]. asymptotics [32]. averaging [63]. avoiding [46].

backward [141]. Barak [150]. based [73]. Basic [10, 54]. behavior [143]. behaviour [18]. Benford [54]. Bernoulli [44]. Bethe [99]. between [105]. bias [103]. bodies [124]. boolean [130]. bootstrapping [40]. Bougerol [69]. bounds [111]. boxes [32]. Branching [36, 47, 93]. bridges [116, 101]. Brownian [31, 53, 48, 21, 29, 15, 16, 87]. BRS [117]. BRS-inequality [117].

calculus [149, 3]. Cauchy [130]. central [5]. chain [22, 5]. chains [38, 106, 2]. channels [45]. chaos [82]. characteristics [97, 136, 114]. Characterizations [78, 55]. Chernoff [111]. circular [62]. classifying [118]. cluster [52]. Coagulation [95]. combinatorial [51]. Combinatorics [52]. comparison [22, 92]. complete [85]. completely [65]. Compound [128]. condensation [64]. conditioned [64]. Conditions [10, 104]. Conductance [58]. Conformal [11, 87]. Conformally [56]. conjectures [148]. constants [31]. constructions [29]. continuous [47, 24, 4]. continuous-state [47]. continuum [134, 37]. control [19, 131]. Controlled [13]. convex [124]. Convolutions [42, 130, 125]. Copulas [96]. cosmology [134]. counterexamples [105]. coupling [71]. critical [56]. crossings [26]. Crump [93]. cumulants [127]. Current [85]. cylindrical [145].

definitions [105]. deletion [45]. delocalization [23]. dependence [40]. descendants [66]. Determinantal [25]. developments [7]. deviations [3, 105, 129]. Differentiability [149]. Differential [38, 1, 141]. diffusion [13, 95, 116, 16]. diffusions [36]. digits [79]. dimension [149]. dimensional [86, 89, 116]. directed [150]. Dirichlet [42]. discrete [49, 60, 70, 112]. discrete-time [49]. Distribution [79, 100]. distributions [90, 92, 68, 132]. divisible [120]. driven [86]. duality [76]. dynamics [152, 94].

effects [130]. eigenvalues [91]. Elementary [145]. embedding [6]. ensembles [17]. enumeration [31]. equation [152, 38, 84]. equations [1, 141, 138, 136, 7]. Equidistribution [100]. Equivalences [105]. erased [98]. Erdos [150]. Erratum [70]. error [143]. essay [28]. everything [84]. evolution [72, 61, 84]. evolutions [129]. exact [137]. examples [42]. exchange [94]. Exchangeable [9, 39, 115]. excursion [31]. Existence [37]. expansions [52]. explain [84]. explicit [42]. Exponential [12, 15, 16]. exposition [99]. expression [108]. extendibility [115]. extensions [50, 69]. extreme [91].

families [130]. FCLT [35]. fields [91, 147, 88, 134]. finance [119]. finite [91, 8]. finite-volume [91]. fixed [15]. Floating [124]. flows [4]. fluctuation [137]. formulas [137]. Fractional [88]. fragmentation [138]. framework [146]. free [130]. Fringe [93]. function [79, 77]. Functional [46, 126]. functionals [12, 122, 26, 15, 16]. functions [65]. Fundamentals [151, 57].

Galton [64]. Gamma [42, 53, 48]. Gaussian [147, 26, 88, 82, 104]. gene [108]. General [2, 32, 28]. Generalized [42]. generalizing [146]. generated [64]. Geometry [131, 60, 70, 21]. GIG [78]. glimpse [72]. graph [144, 31]. graphs [39, 150]. Gromov [146]. groups [8].

Hamiltonian [91]. Hausdorff [146]. heavy [34]. heavy-traffic [34]. Hidden [81]. homogeneous [135]. Horton [109]. hydrodynamics [97]. Hyperbolic [89]. hypergraphs [39]. hypoelliptic [136].

i.i.d [91]. identity [69]. II [16]. impact [83]. income [43]. incompressible [7]. Independence [25]. inequalities [140, 71]. inequality [117]. Infinite [86, 125, 149, 89, 115]. infinitely [32, 120]. information [131]. integers [135, 79]. Integrable [75]. Integral [145, 46]. integrals [18]. interfaces [23]. interval [50]. introduction [152]. invariance [20]. invariant [56]. investment [43]. Itô [145]. Ito [119].

Jagers [93]. jumps [1, 81].

Kernel [130]. KPZ [133].

Lamperti [47]. Large [3, 129, 39, 105]. Last [150, 123]. lattice [37]. law [62, 69, 54]. laws [32, 78, 109, 15, 120]. lecture [63]. Level [26]. Lévy [86, 12, 119]. like [112]. Limit [49, 122, 150, 5, 37, 74, 104]. Limits [112, 101, 61, 34, 56]. linear [136]. Local [114]. Localization [23]. Loewner [72, 61, 129]. long [96]. loop [98]. loop-erased [98].

Macdonald [75]. Malliavin [149]. many [32, 34]. many-server [34]. maps [50]. Markov [38, 22, 65, 76, 5, 106, 2, 142]. Markovian [34]. Martingale [34, 35]. martingales [114]. Mathematical [108, 109]. matrices [83, 3]. matrix [67, 139, 137]. Maximal [140]. maximization [33]. maximum [116]. may [84]. MCMC [2]. meanders [116]. means [42]. measure [24, 80]. measure-theoretical [80]. measures [89, 42, 98, 81, 131, 125]. media [36]. Meixner [55]. memory [96]. metapopulation [49]. Metastable [106]. method [127, 136, 92, 57, 132]. methods [73, 141, 139]. metric [146, 81]. microscopic [97]. mixability [85]. mixed [90]. Mixing [10]. Mode [93]. Model [58]. Models [144, 119, 49, 59, 99, 150, 102, 43, 108]. moment [90]. Moments [53, 48]. monotone [65]. motion [29, 15, 16, 87]. multiple [18]. multiplicative [82]. Multivariate [67].

Navier [7]. Necessary [104]. networks [41]. noise [120]. non [27]. non-uniqueness [27]. Nonclassical [4]. nonnegative [111]. normal [127]. normalized [74, 33]. Notes [32]. notion [40, 76]. Numerical [141, 120].

occupancy [32]. offspring [6]. one [103, 116]. one-dimensional [116]. Open [10, 85]. operators [83]. Opinion [94]. order [136]. Ornstein [86]. Orthogonal [17]. oscillating [135]. other [31, 26]. overview [7]. pairs [9]. partial [151]. particle [37]. passage [123, 150]. PDE [95]. percolation [72, 147, 123, 150, 27, 56]. perspective [95, 100]. physics [134]. place [116]. Planar [72, 56]. point [80]. Poisson [128, 9, 90, 107, 121]. Polchinski [152]. Polish [125]. polynomial [17]. polytopes [124]. population [68]. positive [24]. power [32]. predicting [118]. Prediction [126, 143, 67]. principle [105]. principles [20]. probabilist [100]. Probabilistic [138, 66, 95]. Probability [15, 7, 75, 83, 17, 125]. problem [83, 32, 115, 6]. problems [59, 110]. process [63, 145, 53, 48, 55, 84]. Processes [25, 86, 12, 13, 75, 47, 145, 122, 65, 93, 76, 26, 80, 16, 68, 118, 28, 30, 44, 142, 120, 33]. products [4]. progress [58]. Proof [47, 139]. Proofs [35, 34]. Properties [10, 77]. Pseudo [33]. Pseudo-maximization [33].

quadratic [122, 104]. Quantile [71]. Quasi [68]. Quasi-stationary [68]. Questions [10, 85, 11]. queueing [41]. queues [34].

Random [109, 14, 113, 91, 39, 135, 79, 36, 24, 139, 137, 150, 144, 51, 60, 70, 3, 8, 64, 148, 115, 18, 134, 30, 112, 23, 58]. real [115]. real-valued [115]. realization [24]. Reciprocal [80]. Regeneration [51]. regular [81]. Regularly [81]. reinforcement [30]. rejection [151]. related [59, 150, 16, 45, 11, 120]. representation [75, 47, 120]. representations [46, 138]. restriction [11, 87]. results [137, 8, 45]. review [137, 82]. Reviewing [55]. RSK [123]. Ruin [43].

s [47, 76]. sampling [151]. Sandpile [102]. Scaling [61, 101, 56]. Schramm [72, 61, 129]. search [93]. second [136]. second-order [136]. selection [110]. Self [74, 46, 109, 33]. self-avoiding [46]. Self-normalized [74, 33]. self-similar [109]. semigroups [125]. semimartingales [142]. sensitivity [73]. sequences [143, 90, 20, 104]. sequential [110]. series [126, 40]. server [34]. several [105]. shot [120]. similar [109]. Simply [64]. six [99]. six-vertex [99]. Size [103]. skew [29]. Skorokhod [6]. Smoluchowski [95]. Smooth [147, 50]. solving [110]. Some [10, 16, 133, 140, 19, 7]. space [24, 2]. spaces [89, 81]. spatial [37]. spectral [73]. spin [59]. spinor [134]. spinor-valued [134]. Stability [41]. stable [132]. state [47, 2]. stationary [143, 126, 122, 26, 68, 20, 118]. Statistical [77]. statistics [39, 18]. Stein [92, 57, 132]. Stieltjes [130]. Stochastic [1, 152, 44, 141, 136, 3, 28, 19, 131, 4]. Stokes [7]. strategic [131]. Strong [10]. structures [51, 131, 112, 133]. subtrees [144]. sufficient [104]. sum [79]. sum-of-digits [79]. superdiffusions [36]. supermartingales [111]. supremum [53, 48]. surfaces [21]. Survey [10, 79, 141, 130, 8, 78, 88, 45, 30, 74, 129]. Symbolic [50]. synchronization [45]. systems [37]. Szego [67, 66].

Tail [18]. tangency [114]. TASEP [97]. temporally [65]. Tensor [134]. Tensor- [134]. their [71, 19]. theorem [66, 5]. theorems [49, 150, 122, 60, 70, 37, 74, 104]. theoretical [80]. theory [54, 67, 126, 75, 139, 137, 27, 17, 109, 28]. Thorin [42]. Three [60, 70]. Time [111, 116, 126, 49, 40, 15]. Time-uniform [111]. Toeplitz [83, 122]. Toeplitz-type [122]. Topics [59, 98]. topologies [131]. trace [83]. traffic [34]. transformations [24]. tree [112]. tree-like [112]. trees [93, 64, 109, 14]. Tsirelson [84]. type [122, 53, 48].

Uhlenbeck [86]. unified [123, 110, 146, 132]. uniform [111, 100, 105]. Uniqueness [27]. univariate [92]. universal [118]. Universality [148, 133]. using [97].

valued [115, 134, 114]. values [91]. variables [39, 24]. variance [73]. variation [81]. variations [104]. varying [81]. vector [114]. vector-valued [114]. vectors [115]. vertex [99]. via [24, 111]. view [80]. volume [91].

walk [46, 98, 148]. Walks [113, 135, 8]. Watson [64]. weak [40]. weakly [101]. Wiener [145, 24]. Wright [31].

XXZ [99].

zeros [77]. zeta [77].

References

Bass:2004:SDE

[1] Richard F. Bass. Stochastic differential equations with jumps. Probability Surveys, 1(??):1–19, 2004. CODEN ???? ISSN 1549-5787. URL http://

     projecteuclid.org/euclid.ps/1099928647.

Roberts:2004:GSS

[2] Gareth O. Roberts and Jeffrey S. Rosenthal. General state space Markov chains and MCMC algorithms. Probability Surveys, 1(??):20–71, 2004. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.org/

     euclid.ps/1099928648.

Guionnet:2004:LDS

[3] Alice Guionnet. Large deviations and stochastic calculus for large random matrices. Probability Surveys, 1(??):72–172, 2004. CODEN ???? ISSN

     1549-5787. URL http://projecteuclid.org/euclid.ps/1099928649.

Tsirelson:2004:NSF

[4] Boris Tsirelson. Nonclassical stochastic flows and continuous products. Probability Surveys, 1(??):173–298, 2004. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1104335300.

Jones:2004:MCC

[5] Galin L. Jones. On the Markov chain central limit theorem. Probability Surveys, 1(??):299–320, 2004. CODEN ???? ISSN 1549-5787. URL http:

//projecteuclid.org/euclid.ps/1104335301.

Obloj:2004:SEP

[6] Jan Obłój. The Skorokhod embedding problem and its offspring. Probability Surveys, 1(??):321–392, 2004. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1104335302.

Waymire:2005:PIN

[7] Edward C. Waymire. Probability & incompressible Navier–Stokes equations: An overview of some recent developments. Probability Surveys, 2(??):1–32, 2005. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1109608866.

Hildebrand:2005:SRR

[8] Martin Hildebrand. A survey of results on random random walks on finite groups. Probability Surveys, 2(??):33–63, 2005. CODEN ???? ISSN 1549-

5787. URL http://projecteuclid.org/euclid.ps/1109608867.

Chatterjee:2005:EPP

[9] Sourav Chatterjee, Persi Diaconis, and Elizabeth Meckes. Exchangeable pairs and Poisson approximation. Probability Surveys, 2(??):64–106, 2005. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.org/

euclid.ps/1109608868.

Bradley:2005:BPS

[10] Richard C. Bradley. Basic properties of strong mixing conditions. A survey and some open questions. Probability Surveys, 2(??):107–144, 2005. CO-

DEN ???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.
ps/1115386870.

Werner:2005:CRR

[11] Wendelin Werner. Conformal restriction and related questions. Probability Surveys, 2(??):145–190, 2005. CODEN ???? ISSN 1549-5787. URL http:

//projecteuclid.org/euclid.ps/1117553605.

Bertoin:2005:EFL

[12] Jean Bertoin and Marc Yor. Exponential functionals of Lévy processes. Probability Surveys, 2(??):191–212, 2005. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1127136329.

Borkar:2005:CDP

[13] Vivek S. Borkar. Controlled diffusion processes. Probability Surveys, 2 (??):213–244, 2005. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1127136330.

LeGall:2005:RTA

[14] Jean-François Le Gall. Random trees and applications. Probability Surveys, 2(??):245–311, 2005. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1132583290.

Matsumoto:2005:EFBa

[15] Hiroyuki Matsumoto and Marc Yor. Exponential functionals of Brownian motion, I: Probability laws at fixed time. Probability Surveys, 2(??):312– 347, 2005. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.

org/euclid.ps/1132583291.

Matsumoto:2005:EFBb

[16] Hiroyuki Matsumoto and Marc Yor. Exponential functionals of Brownian motion, II: Some related diffusion processes. Probability Surveys, 2(??):348–384, 2005. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1132583292.

Konig:2005:OPE

[17] Wolfgang König. Orthogonal polynomial ensembles in probability theory. Probability Surveys, 2(??):385–447, 2005. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1133360670.

Major:2005:TBM

[18] Péter Major. Tail behaviour of multiple random integrals and U-statistics. Probability Surveys, 2(??):448–505, 2005. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1134397603.

Pham:2005:SRA

[19] Huyên Pham. On some recent aspects of stochastic control and their applications. Probability Surveys, 2(??):506–549, 2005. CODEN ???? ISSN

1549-5787. URL http://projecteuclid.org/euclid.ps/1135911648.

Merlevede:2006:RAI

[20] Florence Merlevède, Magda Peligrad, and Sergey Utev. Recent advances in invariance principles for stationary sequences. Probability Surveys, 3(??):1–36, 2006. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1137162933.

Leandre:2006:GBS

[21] Rémi Léandre. The geometry of Brownian surfaces. Probability Surveys, 3(??):37–88, 2006. CODEN ???? ISSN 1549-5787. URL http:/

/projecteuclid.org/euclid.ps/1145452120.

Dyer:2006:MCC

[22] Martin Dyer, Leslie Ann Goldberg, Mark Jerrum, and Russell Martin. Markov chain comparison. Probability Surveys, 3(??):89–111, 2006. CO-

DEN ???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.
ps/1145890796.

Velenik:2006:LDR

[23] Yvan Velenik. Localization and delocalization of random interfaces. Probability Surveys, 3(??):112–169, 2006. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1146757332.

Feyel:2006:RPR

[24] D. Feyel, A. S. Üstünel, and M. Zakai. The realization of positive random variables via absolutely continuous transformations of measure on Wiener space. Probability Surveys, 3(??):170–205, 2006. CODEN ???? ISSN

1549-5787. URL http://projecteuclid.org/euclid.ps/1146832695.

Hough:2006:DPI

[25] J. Ben Hough, Manjunath Krishnapur, Yuval Peres, and Bálint Virág. Determinantal processes and independence. Probability Surveys, 3(??):206– 229, 2006. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.

org/euclid.ps/1146832696.

Kratz:2006:LCO

[26] Marie F. Kratz. Level crossings and other level functionals of stationary Gaussian processes. Probability Surveys, 3(??):230–288, 2006. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1166541148.

Haggstrom:2006:UNU

[27] Olle Häggström and Johan Jonasson. Uniqueness and non-uniqueness in percolation theory. Probability Surveys, 3(??):289–344, 2006. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1167536197.

Nikeghbali:2006:EGT

[28] Ashkan Nikeghbali. An essay on the general theory of stochastic processes. Probability Surveys, 3(??):345–412, 2006. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1167752810.

Lejay:2006:CSB

[29] Antoine Lejay. On the constructions of the skew Brownian motion. Probability Surveys, 3(??):413–466, 2006. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1168268869.

Pemantle:2007:SRP

[30] Robin Pemantle. A survey of random processes with reinforcement. Probability Surveys, 4(??):1–79, 2007. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1172244137.

Janson:2007:BEA

[31] Svante Janson. Brownian excursion area, Wright’s constants in graph enumeration, and other Brownian areas. Probability Surveys, 4(??):80– 145, 2007. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.

org/euclid.ps/1178804352.

Gnedin:2007:NOP

[32] Alexander Gnedin, Ben Hansen, and Jim Pitman. Notes on the occupancy problem with infinitely many boxes: general asymptotics and power laws. Probability Surveys, 4(??):146–171, 2007. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1180728778.

delaPena:2007:PMS

[33] Victor H. de la Peña, Michael J. Klass, and Tze Leung Lai. Pseudomaximization and self-normalized processes. Probability Surveys, 4(??): 172–192, 2007. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1192118046.

Pang:2007:MPM

[34] Guodong Pang, Rishi Talreja, and Ward Whitt. Martingale proofs of many-server heavy-traffic limits for Markovian queues. Probability Surveys, 4(??):193–267, 2007. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1200511983.

Whitt:2007:PMF

[35] Ward Whitt. Proofs of the martingale FCLT. Probability Surveys, 4 (??):268–302, 2007. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1198788218.

Englander:2007:BDS

[36] János Engländer. Branching diffusions, superdiffusions and random media. Probability Surveys, 4(??):303–364, 2007. CODEN ???? ISSN 1549-

5787. URL http://projecteuclid.org/euclid.ps/1200511984.

Penrose:2008:ESL

[37] Mathew D. Penrose. Existence and spatial limit theorems for lattice and continuum particle systems. Probability Surveys, 5(??):1–36, 2008. CO-

DEN ???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.
ps/1207254889.

Darling:2008:DEA

[38] R. W. R. Darling and J. R. Norris. Differential equation approximations for Markov chains. Probability Surveys, 5(??):37–79, 2008. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1208958281.

Austin:2008:ERV

[39] Tim Austin. On exchangeable random variables and the statistics of large graphs and hypergraphs. Probability Surveys, 5(??):80–145, 2008. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1211574613.

Doukhan:2008:NWD

[40] Paul Doukhan and Michael H. Neumann. The notion of ψ-weak dependence and its applications to bootstrapping time series. Probability Surveys, 5(??):146–168, 2008. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1214491946.

Bramson:2008:SQN

[41] Maury Bramson. Stability of queueing networks. Probability Surveys, 5(??):169–345, 2008. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1220879338.

James:2008:GGC

[42] Lancelot F. James, Bernard Roynette, and Marc Yor. Generalized gamma convolutions, Dirichlet means, Thorin measures, with explicit examples.

Probability Surveys, 5(??):346–415, 2008. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1223654264.

Paulsen:2008:RMI

[43] Jostein Paulsen. Ruin models with investment income. Probability Surveys, 5(??):416–434, 2008. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1229524758.

Privault:2008:SAB

[44] Nicolas Privault. Stochastic analysis of Bernoulli processes. Probability Surveys, 5(??):435–483, 2008. CODEN ???? ISSN 1549-5787. URL http:

//projecteuclid.org/euclid.ps/1230559281.

Mitzenmacher:2009:SRD

[45] Michael Mitzenmacher. A survey of results for deletion channels and related synchronization channels. Probability Surveys, 6(??):1–33, 2009. CO-

DEN ???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.
ps/1247144661.

Brydges:2009:FIR

[46] David C. Brydges, John Z. Imbrie, and Gordon Slade. Functional integral representations for self-avoiding walk. Probability Surveys, 6(??):34–61, 2009. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.

org/euclid.ps/1249996530.

Caballero:2009:PLR

[47] Ma. Emilia Caballero, Amaury Lambert, and Gerónimo Uribe Bravo. Proof(s) of the Lamperti representation of continuous-state branching processes. Probability Surveys, 6(??):62–89, 2009. CODEN ???? ISSN 1549-

5787. URL http://projecteuclid.org/euclid.ps/1259677176.

Janson:2010:MGT

[48] Svante Janson. Moments of Gamma type and the Brownian supremum process area. Probability Surveys, 7(??):1–52, 2010. CODEN ???? ISSN

1549-5787. URL http://projecteuclid.org/euclid.ps/1272029738.

See addendum [53].

Buckley:2010:LTD

[49] F. M. Buckley and P. K. Pollett. Limit theorems for discrete-time metapopulation models. Probability Surveys, 7(??):53–83, 2010. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1273670365.

Downarowicz:2010:SES

[50] Tomasz Downarowicz. Symbolic extensions of smooth interval maps. Probability Surveys, 7(??):84–104, 2010. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1274198273.

Gnedin:2010:RRC

[51] Alexander V. Gnedin. Regeneration in random combinatorial structures. Probability Surveys, 7(??):105–156, 2010. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1274198274.

Faris:2010:CCE

[52] William G. Faris. Combinatorics and cluster expansions. Probability Surveys, 7(??):157–206, 2010. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1275928347.

Janson:2010:AMG

[53] Svante Janson. Addendum to moments of Gamma type and the Brownian supremum process area. Probability Surveys, 7(??):207–208, 2010. CO-

DEN ???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.
ps/1292249775. See [48].

Berger:2011:BTB

[54] Arno Berger and Theodore P. Hill. A basic theory of Benford’s Law. Probability Surveys, 8(??):1–126, 2011. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1311860830.

Mazzola:2011:RAC

[55] E. Mazzola and P. Muliere. Reviewing alternative characterizations of Meixner process. Probability Surveys, 8(??):127–154, 2011. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1311860831.

Sun:2011:CIS

[56] Nike Sun. Conformally invariant scaling limits in planar critical percolation. Probability Surveys, 8(??):155–209, 2011. CODEN ???? ISSN

1549-5787. URL http://projecteuclid.org/euclid.ps/1319806861.

Ross:2011:FSM

[57] Nathan Ross. Fundamentals of Stein’s method. Probability Surveys, 8 (??):210–293, 2011. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1319806862.

Biskup:2011:RPR

[58] Marek Biskup. Recent progress on the Random Conductance Model. Probability Surveys, 8(??):294–373, 2011. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1325264815.

Dubedat:2011:TAS

[59] Julien Dubédat. Topics on abelian spin models and related problems. Probability Surveys, 8(??):374–402, 2011. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1325264816.

Grimmett:2011:TTD

[60] Geoffrey Grimmett. Three theorems in discrete random geometry. Probability Surveys, 8(??):403–441, 2011. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1325264817. See erratum [70].

Lawler:2011:SLS

[61] Gregory F. Lawler. Scaling limits and the Schramm–Loewner evolution. Probability Surveys, 8(??):442–495, 2011. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1325264818.

Bordenave:2012:ACL

[62] Charles Bordenave and Djalil Chafaï. Around the circular law. Probability Surveys, 9(??):1–89, 2012. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1325604980.

Aldous:2012:LAP

[63] David Aldous and Daniel Lanoue. A lecture on the averaging process. Probability Surveys, 9(??):90–102, 2012. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1327328305.

Janson:2012:SGT

[64] Svante Janson. Simply generated trees, conditioned Galton–Watson trees, random allocations and condensation. Probability Surveys, 9(??):103–252, 2012. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.org/

euclid.ps/1331216239.

Hirsch:2012:TCM

[65] Francis Hirsch and Marc Yor. On temporally completely monotone functions for Markov processes. Probability Surveys, 9(??):253–286, 2012. CO-

DEN ???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.
ps/1336658304.

Bingham:2012:STP

[66] N. H. Bingham. Szegő’s theorem and its probabilistic descendants. Probability Surveys, 9(??):287–324, 2012. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1343047754.

Bingham:2012:MPM

[67] N. H. Bingham. Multivariate prediction and matrix Szegő theory. Probability Surveys, 9(??):325–339, 2012. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1343047755.

Meleard:2012:QSD

[68] Sylvie Méléard and Denis Villemonais. Quasi-stationary distributions and population processes. Probability Surveys, 9(??):340–410, 2012. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1350047379.

Vakeroudis:2012:BIL

[69] Stavros Vakeroudis. Bougerol’s identity in law and extensions. Probability Surveys, 9(??):411–437, 2012. CODEN ???? ISSN 1549-5787. URL http:

//projecteuclid.org/euclid.ps/1352385533.

Grimmett:2012:ETT

[70] Geoffrey Grimmett. Erratum: Three theorems in discrete random geometry. Probability Surveys, 9(??):438, 2012. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1352903647. See [60].

Mason:2012:QCI

[71] David M. Mason and Harrison H. Zhou. Quantile coupling inequalities and their applications. Probability Surveys, 9(??):439–479, 2012. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1354125785.

Beffara:2013:PPG

[72] Vincent Beffara and Hugo Duminil-Copin. Planar percolation with a glimpse of Schramm–Loewner evolution. Probability Surveys, 10(??):1–50, 2013. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.org/

euclid.ps/1379686423.

Alexanderian:2013:SMV

[73] Alen Alexanderian. On spectral methods for variance based sensitivity analysis. Probability Surveys, 10(??):51–68, 2013. CODEN ???? ISSN

1549-5787. URL http://projecteuclid.org/euclid.ps/1385129850.

Shao:2013:SNL

[74] Qi-Man Shao and Qiying Wang. Self-normalized limit theorems: A survey. Probability Surveys, 10(??):69–93, 2013. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1385665279.

Borodin:2014:IPR

[75] Alexei Borodin and Leonid Petrov. Integrable probability: From representation theory to Macdonald processes. Probability Surveys, 11(??):1–58, 2014. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.org/

euclid.ps/1395076923.

Jansen:2014:NDM

[76] Sabine Jansen and Noemi Kurt. On the notion(s) of duality for Markov processes. Probability Surveys, 11(??):59–120, 2014. CODEN ???? ISSN

1549-5787. URL http://projecteuclid.org/euclid.ps/1398778562.

Kargin:2014:SPZ

[77] Vladislav Kargin. Statistical properties of zeta functions’ zeros. Probability Surveys, 11(??):121–160, 2014. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1404393850.

Koudou:2014:CGL

[78] Angelo Efoévi Koudou and Christophe Ley. Characterizations of GIG laws: A survey. Probability Surveys, 11(??):161–176, 2014. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1405603141.

Chen:2014:DSD

[79] Louis H. Y. Chen, Hsien-Kuei Hwang, and Vytas Zacharovas. Distribution of the sum-of-digits function of random integers: A survey. Probability Surveys, 11(??):177–236, 2014. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1412947841.

Leonard:2014:RPM

[80] Christian Léonard, Sylvie Rœlly, and Jean-Claude Zambrini. Reciprocal processes. A measure-theoretical point of view. Probability Surveys, 11(??):237–269, 2014. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1412947842.

Lindskog:2014:RVM

[81] Filip Lindskog, Sidney I. Resnick, and Joyjit Roy. Regularly varying measures on metric spaces: Hidden regular variation and hidden jumps. Probability Surveys, 11(??):270–314, 2014. CODEN ???? ISSN 1549-5787.

URL http://projecteuclid.org/euclid.ps/1413896892.

Rhodes:2014:GMC

[82] Rémi Rhodes and Vincent Vargas. Gaussian multiplicative chaos and applications: A review. Probability Surveys, 11(??):315–392, 2014. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1415023603.

Ginovyan:2014:TPT

[83] Mamikon S. Ginovyan, Artur A. Sahakyan, and Murad S. Taqqu. The trace problem for Toeplitz matrices and operators and its impact in probability. Probability Surveys, 11(??):393–440, 2014. CODEN ???? ISSN

1549-5787. URL http://projecteuclid.org/euclid.ps/1417528609.

Yano:2015:ATE

[84] Kouji Yano and Marc Yor. Around Tsirelson’s equation, or: The evolution process may not explain everything. Probability Surveys, 12(??):1–12, 2015. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.

org/euclid.ps/1437497795.

Wang:2015:COQ

[85] Ruodu Wang. Current open questions in complete mixability. Probability Surveys, 12(??):13–32, 2015. CODEN ???? ISSN 1549-5787. URL http:

//projecteuclid.org/euclid.ps/1440075824.

Applebaum:2015:IDO

[86] David Applebaum. Infinite dimensional Ornstein–Uhlenbeck processes driven by Lévy processes. Probability Surveys, 12(??):33–54, 2015. CO-

DEN ???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.
ps/1440075825.

Wu:2015:CRB

[87] Hao Wu. Conformal restriction and Brownian motion. Probability Surveys, 12(??):55–103, 2015. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1444653628.

Lodhia:2016:FGF

[88] Asad Lodhia, Scott Sheffield, Xin Sun, and Samuel S. Watson. Fractional Gaussian fields: A survey. Probability Surveys, 13(??):1–56, 2016. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1456149586.

Bobkov:2016:HMI

[89] Sergey G. Bobkov and James Melbourne. Hyperbolic measures on infinite dimensional spaces. Probability Surveys, 13(??):57–88, 2016. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1465321421.

Kuba:2016:MSM

[90] Markus Kuba and Alois Panholzer. On moment sequences and mixed Poisson distributions. Probability Surveys, 13(??):89–155, 2016. CODEN

???? ISSN 1549-5787. URL http://projecteuclid.org/euclid.ps/
1474374818.

Astrauskas:2016:EVD

[91] Arvydas Astrauskas. From extreme values of i.i.d. random fields to extreme eigenvalues of finite-volume Anderson Hamiltonian. Probability Surveys, 13(??):156–244, 2016. CODEN ???? ISSN 1549-5787. URL http:/

/projecteuclid.org/euclid.ps/1476369041.

Ley:2017:SMC

[92] Christophe Ley, Gesine Reinert, and Yvik Swan. Stein’s method for comparison of univariate distributions. Probability Surveys, 14(??):1–52, 2017. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.org/

euclid.ps/1483952471.

Holmgren:2017:FTC

[93] Cecilia Holmgren and Svante Janso. Fringe trees, Crump–Mode–Jagers branching processes and m-ary search trees. Probability Surveys, 14 (??):53–154, 2017. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1490169611.

Mossel:2017:OED

[94] Elchanan Mossel and Omer Tamuz. Opinion exchange dynamics. Probability Surveys, 14:155–204, 2017. CODEN ???? ISSN 1549-5787.

Hammond:2017:CDP

[95] Alan Hammond. Coagulation and diffusion: A probabilistic perspective on the Smoluchowski PDE. Probability Surveys, 14:205–288, 2017. CODEN ???? ISSN 1549-5787.

Ibragimov:2017:CLM

[96] Rustam Ibragimov and George Lentzas. Copulas and long memory. Probability Surveys, 14:289–327, 2017. CODEN ???? ISSN 1549-5787.

Ferrari:2018:THU

[97] Pablo A. Ferrari. TASEP hydrodynamics using microscopic characteristics. Probability Surveys, 15:1–27, 2018. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/euclid.ps/1519722018.

Lawler:2018:TLM

[98] Gregory F. Lawler. Topics in loop measures and the loop-erased walk. Probability Surveys, 15(??):28–101, 2018. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/euclid.ps/1520326908.

Duminil-Copin:2018:BAS

[99] Hugo Duminil-Copin, Maxime Gagnebin, Matan Harel, Ioan Manolescu, and Vincent Tassion. The Bethe ansatz for the six-vertex and XXZ models: An exposition. Probability Surveys, 15(??):102–130, 2018. CODEN

???? ISSN 1549-5787. URL https://projecteuclid.org/euclid.ps/
1521079210.

Limic:2018:EUD

[100] Vlada Limic and Nedzad Limić. Equidistribution, uniform distribution: a probabilist’s perspective. Probability Surveys, 15(??):131–155, 2018. CO-

DEN ???? ISSN 1549-5787. URL https://projecteuclid.org/euclid.
ps/1524556822.

Labbe:2018:SLW

[101] Cyril Labbé. On the scaling limits of weakly asymmetric bridges. Probability Surveys, 15(??):156–242, 2018. CODEN ???? ISSN 1549-5787. URL

https://projecteuclid.org/euclid.ps/1537408916.

Jarai:2018:SM

[102] Antal A. Járai. Sandpile models. Probability Surveys, 15(??):243–306, 2018. CODEN ???? ISSN 1549-5787. URL https://projecteuclid.

org/euclid.ps/1537776024.

Arratia:2019:SBO

[103] Richard Arratia, Larry Goldstein, and Fred Kochman. Size bias for one and all. Probability Surveys, 16(??):1–61, 2019. CODEN ???? ISSN

1549-5787. URL https://projecteuclid.org/euclid.ps/1546657438.

Viitasaari:2019:NSC

[104] Lauri Viitasaari. Necessary and sufficient conditions for limit theorems for quadratic variations of Gaussian sequences. Probability Surveys, 16 (??):62–98, 2019. CODEN ???? ISSN 1549-5787. URL https://

projecteuclid.org/euclid.ps/1559289658.

Salins:2019:ECB

[105] Michael Salins. Equivalences and counterexamples between several definitions of the uniform large deviations principle. Probability Surveys, 16(??):99–142, 2019. CODEN ???? ISSN 1549-5787. URL https://

projecteuclid.org/euclid.ps/1559289659.

Landim:2019:MMC

[106] Claudio Landim. Metastable Markov chains. Probability Surveys, 16 (??):143–227, 2019. CODEN ???? ISSN 1549-5787. URL https://

projecteuclid.org/euclid.ps/1560153809.

Novak:2019:PA

[107] S. Y. Novak. Poisson approximation. Probability Surveys, 16(??):228–276, 2019. CODEN ???? ISSN 1549-5787. URL https://projecteuclid.

org/euclid.ps/1565748164.

Robert:2019:MMG

[108] Philippe Robert. Mathematical models of gene expression. Probability Surveys, 16(??):277–332, 2019. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1572509200.

Kovchegov:2020:RSS

[109] Yevgeniy Kovchegov and Ilya Zaliapin. Random self-similar trees: A mathematical theory of Horton laws. Probability Surveys, 17(??):1–213, 2020. CODEN ???? ISSN 1549-5787. URL http://projecteuclid.org/

euclid.ps/1582340412.

Goldenshluger:2020:UAS

[110] Alexander Goldenshluger, Yaakov Malinovsky, and Assaf Zeevi. A unified approach for solving sequential selection problems. Probability Surveys, 17(??):214–256, 2020. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1587974426.

Howard:2020:TUC

[111] Steven R. Howard, Aaditya Ramdas, Jon McAuliffe, and Jasjeet Sekhon. Time-uniform Chernoff bounds via nonnegative supermartingales. Probability Surveys, 17(??):257–317, 2020. CODEN ???? ISSN 1549-5787. URL

http://projecteuclid.org/euclid.ps/1589940218.

Stufler:2020:LRT

[112] Benedikt Stufler. Limits of random tree-like discrete structures. Probability Surveys, 17(??):318–477, 2020. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-17/issue-none/Limits-of-random-tree-like-discrete-structures/
10.1214/19-PS338.full.

Rolla:2020:ARW

[113] Leonardo T. Rolla. Activated random walks on Z^{d}. Probability Surveys, 17(??):478–544, 2020. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-17/issue-none/Activated-Random-Walks-on-mathbbZd/10.1214/
19-PS339.full.

Yaroslavtsev:2020:LCT

[114] Ivan S. Yaroslavtsev. Local characteristics and tangency of vectorvalued martingales. Probability Surveys, 17(??):545–676, 2020. CO-DEN ???? ISSN 1549-5787. URL https://projecteuclid.

org/journals/probability-surveys/volume-17/issue-none/Local-
characteristics-and-tangency-of-vector-valued-martingales/10.
1214/19-PS337.full.

Mai:2020:IEP

[115] Jan-Frederik Mai. The infinite extendibility problem for exchangeable real-valued random vectors. Probability Surveys, 17(??):677–753, 2020.

CODEN ???? ISSN 1549-5787. URL https://projecteuclid.org/
journals/probability-surveys/volume-17/issue-none/The-infinite-
extendibility-problem-for-exchangeable-real-valued-random-vectors/
10.1214/19-PS336.full.

Khanfir:2021:TPM

[116] Robin Khanfir. Time and place of the maximum for one-dimensional diffusion bridges and meanders. Probability Surveys, 18(??):1–43, 2021.

CODEN ???? ISSN 1549-5787. URL https://projecteuclid.org/
journals/probability-surveys/volume-18/issue-none/Time-and-place-
of-the-maximum-for-one-dimensional-diffusion/10.1214/18-PS312.
full.

Bruss:2021:BIA

[117] F. Thomas Bruss. The BRS-inequality and its applications. Probability Surveys, 18(??):44–76, 2021. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-18/issue-none/The-BRS-inequality-and-its-applications/
10.1214/20-PS351.full.

Morvai:2021:UAC

[118] Gusztáv Morvai and Benjamin Weiss. On universal algorithms for classifying and predicting stationary processes. Probability Surveys, 18 (??):77–131, 2021. CODEN ???? ISSN 1549-5787. URL https://

projecteuclid.org/journals/probability-surveys/volume-18/issue-
none/On-universal-algorithms-for-classifying-and-predicting-
stationary-processes/10.1214/20-PS345.full.

Bouzianis:2021:LIM

[119] George Bouzianis, Lane P. Hughston, Sebastian Jaimungal, and Leandro Sánchez-Betancourt. Lévy–Ito models in finance. Probability Surveys, 18(??):132–178, 2021. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-18/issue-none/L%c3%a9vy-Ito-models-in-finance/10.1214/
21-PS1.full.

Yuan:2021:NAS

[120] Sida Yuan and Reiichiro Kawai. Numerical aspects of shot noise representation of infinitely divisible laws and related processes. Probability Surveys, 18(??):201–271, 2021. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-18/issue-none/Numerical-aspects-of-shot-noise-representation-
of-infinitely-divisible-laws/10.1214/20-PS359.full.

Novak:2021:PAA

[121] S. Y. Novak. Poisson approximation. addendum. Probability Surveys, 18(??):272–275, 2021. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-18/issue-none/Poisson-approximation-Addendum/10.1214/
21-PS2.full.

Ginovyan:2022:LTT

[122] Mamikon S. Ginovyan and Murad S. Taqqu. Limit theorems for Toeplitz-type quadratic functionals of stationary processes and

applications.  Probability Surveys, 19(??):1–64, 2022.  CODEN

???? ISSN 1549-5787. URL https://projecteuclid.org/journals/
probability-surveys/volume-19/issue-none/Limit-theorems-for-
Toeplitz-type-quadratic-functionals-of-stationary-processes/
10.1214/21-PS3.full.

Dauvergne:2022:RLP

[123] Duncan Dauvergne, Mihai Nica, and Bálint Virág. RSK in last passage percolation: a unified approach. Probability Surveys, 19(??):65–112, 2022.

CODEN ???? ISSN 1549-5787. URL https://projecteuclid.org/
journals/probability-surveys/volume-19/issue-none/RSK-in-last-
passage-percolation-a-unified-approach/10.1214/22-PS4.full.

Werner:2022:FBA

[124] Elisabeth M. Werner. Floating bodies and approximation of convex bodies by polytopes. Probability Surveys, 19(??):113–128, 2022. CODEN

???? ISSN 1549-5787. URL https://projecteuclid.org/journals/
probability-surveys/volume-19/issue-none/Floating-bodies-and-
approximation-of-convex-bodies-by-polytopes/10.1214/22-PS5.full.

Yano:2022:ICP

[125] Kouji Yano. Infinite convolutions of probability measures on Polish semigroups. Probability Surveys, 19(??):129–159, 2022. CODEN ???? ISSN

1549-5787. URL https://projecteuclid.org/journals/probability-
surveys/volume-19/issue-none/Infinite-convolutions-of-probability-
measures-on-Polish-semigroups/10.1214/22-PS6.full.

Bingham:2022:PTS

[126] N. H. Bingham. Prediction theory for stationary functional time series. Probability Surveys, 19(??):160–184, 2022. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-19/issue-none/Prediction-theory-for-stationary-functional-
time-series/10.1214/20-PS360.full.

Doring:2022:MCN

[127] Hanna Döring, Sabine Jansen, and Kristina Schubert. The method of cumulants for the normal approximation. Probability Surveys, 19 (??):185–270, 2022. CODEN ???? ISSN 1549-5787. URL https://

projecteuclid.org/journals/probability-surveys/volume-19/issue-
none/The-method-of-cumulants-for-the-normal-approximation/10.
1214/22-PS7.full.

Cekanavicius:2022:CPA

[128] V. Cekanavicius and S. Y. Novak. Compound Poisson approximation. Probability Surveys, 19(??):271–350, 2022. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-19/issue-none/Compound-Poisson-approximation/10.1214/
22-PS8.full.

Wang:2022:LDS

[129] Yilin Wang. Large deviations of Schramm–Loewner evolutions: a survey. Probability Surveys, 19(??):351–403, 2022. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-19/issue-none/Large-deviations-of-Schramm-Loewner-evolutions-
A-survey/10.1214/22-PS9.full.

Fakhfakh:2022:SEF

[130] Raouf Fakhfakh. A survey on the effects of free and boolean convolutions on Cauchy–Stieltjes kernel families. Probability Surveys, 19(??):404–449, 2022. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-19/issue-none/A-survey-on-the-effects-of-free-and-boolean-
convolutions/10.1214/22-PS10.full.

Saldi:2022:GIS

[131] Naci Saldi and Serdar Yüksel. Geometry of information structures, strategic measures and associated stochastic control topologies. Probability Surveys, 19(??):450–532, 2022. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-19/issue-none/Geometry-of-information-structures-strategic-
measures-and-associated-stochastic-control/10.1214/20-PS356.
full.

Upadhye:2022:UAS

[132] Neelesh S. Upadhye and Kalyan Barman. A unified approach to Stein’s method for stable distributions. Probability Surveys, 19(??):533–589, 2022.

CODEN ???? ISSN 1549-5787. URL https://projecteuclid.org/
journals/probability-surveys/volume-19/issue-none/A-unified-
approach-to-Steins-method-for-stable-distributions/10.1214/20-
PS354.full.

Zygouras:2022:SAS

[133] Nikos Zygouras. Some algebraic structures in KPZ universality. Probability Surveys, 19(??):590–700, 2022. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-19/issue-none/Some-algebraic-structures-in-KPZ-universality/
10.1214/19-PS335.full.

Malyarenko:2023:TSV

[134] Anatoliy Malyarenko and Martin Ostoja-Starzewski. Tensor- and spinor-valued random fields with applications to continuum physics and cosmology. Probability Surveys, 20(??):1–86, 2023. CODEN

???? ISSN 1549-5787. URL https://projecteuclid.org/journals/
probability-surveys/volume-20/issue-none/Tensor--and-spinor-
valued-random-fields-with-applications-to/10.1214/22-PS12.full.

Bremont:2023:HOR

[135] Julien Brémont. On homogeneous and oscillating random walks on the integers. Probability Surveys, 20(??):87–112, 2023. CODEN ???? ISSN

1549-5787. URL https://projecteuclid.org/journals/probability-
surveys/volume-20/issue-none/On-homogeneous-and-oscillating-
random-walks-on-the-integers/10.1214/23-PS13.full.

Foldes:2023:MSC

[136] Juraj Földes and David P. Herzog. The method of stochastic characteristics for linear second-order hypoelliptic equations. Probability Surveys, 20(??):113–169, 2023. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-20/issue-none/The-method-of-stochastic-characteristics-
for-linear-second-order-hypoelliptic/10.1214/22-PS11.full.

Forrester:2023:RER

[137] Peter J. Forrester. A review of exact results for fluctuation formulas in random matrix theory. Probability Surveys, 20(??):170–225, 2023. CODEN

???? ISSN 1549-5787. URL https://projecteuclid.org/journals/
probability-surveys/volume-20/issue-none/A-review-of-exact-results-
for-fluctuation-formulas-in-random/10.1214/23-PS15.full.

Deaconu:2023:PRF

[138] Madalina Deaconu and Antoine Lejay. Probabilistic representations of fragmentation equations. Probability Surveys, 20(??):226–290, 2023.

CODEN ???? ISSN 1549-5787. URL https://projecteuclid.org/
journals/probability-surveys/volume-20/issue-none/Probabilistic-
representations-of-fragmentation-equations/10.1214/23-PS14.full.

Fleermann:2023:PMR

[139] Michael Fleermann and Werner Kirsch. Proof methods in random matrix theory. Probability Surveys, 20(??):291–381, 2023. CODEN ???? ISSN

1549-5787. URL https://projecteuclid.org/journals/probability-
surveys/volume-20/issue-none/Proof-methods-in-random-matrix-
theory/10.1214/23-PS16.full.

Kuhn:2023:MIS

[140] Franziska Kühn and René L. Schilling. Maximal inequalities and some applications. Probability Surveys, 20(??):382–485, 2023. CODEN ???? ISSN

1549-5787. URL https://projecteuclid.org/journals/probability-
surveys/volume-20/issue-none/Maximal-inequalities-and-some-applications/
10.1214/23-PS17.full.

Chessari:2023:NMB

[141] Jared Chessari, Reiichiro Kawai, Yuji Shinozaki, and Toshihiro Yamada. Numerical methods for backward stochastic differential equations: a survey. Probability Surveys, 20(??):486–567, 2023. CODEN ???? ISSN

1549-5787. URL https://projecteuclid.org/journals/probability-
surveys/volume-20/issue-none/Numerical-methods-for-backward-
stochastic-differential-equations-A-survey/10.1214/23-PS18.full.

Schnurr:2023:MPS

[142] Alexander Schnurr and Sebastian Rickelhoff. From Markov processes to semimartingales. Probability Surveys, 20(??):568–607, 2023. CODEN

???? ISSN 1549-5787. URL https://projecteuclid.org/journals/
probability-surveys/volume-20/issue-none/From-Markov-processes-
to-semimartingales/10.1214/23-PS19.full.

Babayan:2023:ABP

[143] Nikolay M. Babayan and Mamikon S. Ginovyan. Asymptotic behavior of the prediction error for stationary sequences. Probability Surveys, 20(??):664–721, 2023. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-20/issue-none/Asymptotic-behavior-of-the-prediction-error-
for-stationary-sequences/10.1214/23-PS21.full.

Fredes:2023:MRS

[144] Luis Fredes and Jean-François Marckert. Models of random subtrees of a graph. Probability Surveys, 20(??):722–801, 2023. CODEN ???? ISSN

1549-5787. URL https://projecteuclid.org/journals/probability-
surveys/volume-20/issue-none/Models-of-random-subtrees-of-a-
graph/10.1214/23-PS22.full.

Dotti:2023:EPI

[145] Sylvain Dotti. Elementary processes for Itô integral against cylindrical Wiener process. Probability Surveys, 20(??):802–836, 2023. CODEN

???? ISSN 1549-5787. URL https://projecteuclid.org/journals/
probability-surveys/volume-20/issue-none/Elementary-processes-
for-It%c3%b4-Integral-against-cylindrical-Wiener-process/10.
1214/23-PS23.full.

Khezeli:2023:UFG

[146] Ali Khezeli. A unified framework for generalizing the Gromov– Hausdorff metric. Probability Surveys, 20(??):837–896, 2023. CO-

DEN ????  ISSN 1549-5787. URL https://projecteuclid.org/

journals/probability-surveys/volume-20/issue-none/A-unified-
framework-for-generalizing-the-Gromov-Hausdorff-metric/10.1214/
20-PS340.full.

Beliaev:2023:SGF

[147] Dmitry Beliaev. Smooth Gaussian fields and percolation. Probability Surveys, 20(??):897–937, 2023. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-20/issue-none/Smooth-Gaussian-fields-and-percolation/
10.1214/23-PS24.full.

Levine:2024:UCA

[148] Lionel Levine and Vittoria Silvestri. Universality conjectures for activated random walk. Probability Surveys, 21(??):1–27, 2024. CODEN ???? ISSN

1549-5787. URL https://projecteuclid.org/journals/probability-
surveys/volume-21/issue-none/Universality-conjectures-for-activated-
random-walk/10.1214/24-PS25.full.

Bignamini:2024:DID

[149] Davide A. Bignamini, Simone Ferrari, Simona Fornaro, and Margherita Zanella. Differentiability in infinite dimension and the Malliavin calculus. Probability Surveys, 21(??):28–66, 2024. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-21/issue-none/Differentiability-in-infinite-dimension-
and-the-Malliavin-calculus/10.1214/24-PS26.full.

Foss:2024:LPP

[150] Sergey Foss, Takis Konstantopoulos, Bastien Mallein, and Sanjay Ramassamy. Last passage percolation and limit theorems in Barak– Erdős directed random graphs and related models. Probability Surveys, 21(??):67–170, 2024. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-21/issue-none/Last-passage-percolation-and-limit-theorems-
in-Barak-Erd%c5%91s-directed/10.1214/24-PS28.full.

Jerrum:2024:FPR

[151] Mark Jerrum. Fundamentals of partial rejection sampling. Probability Surveys, 21(??):171–199, 2024. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-21/issue-none/Fundamentals-of-partial-rejection-sampling/
10.1214/24-PS29.full.

Bauerschmidt:2024:SDP

[152] Roland Bauerschmidt, Thierry Bodineau, and Benoit Dagallier. Stochastic dynamics and the Polchinski equation: an introduction. Probability Surveys, 21(??):200–290, 2024. CODEN ???? ISSN 1549-5787.

URL https://projecteuclid.org/journals/probability-surveys/
volume-21/issue-none/Stochastic-dynamics-and-the-Polchinski-
equation-An-introduction/10.1214/24-PS27.full.